# Jingjing Sun, Pengyi Shi, Jim Dai

### v3

@@ 5.3 Batching Setup and Other Implementation Details
In the general atomic decision process, each of the atomic actions sequentially depends on the atomic states as illustrated in Figure 1. We propose a batching setup to simplify this process, which uses the same initial state $s=s^{1}$ for all atomic actions in a batch. Specifically, given $s$ and a chosen customer class order $\boldsymbol{\sigma}(s)$, the probability of conducting atomic action sequence a becomes $\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))$. In other words, during this batch, since the input state $s$ no longer changes, we prescribe atomic actions to the customers from the same class $i$ following the same distribution $\kappa(\cdot|s,i)$. Under the assumption that any system-level action $f$ satisfying $\sum_{j}f_{i,j}=q_{i},\forall i$ is feasible, we have the following lemma.
@@
$\pi(f|s)=\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$ (14)
@@
$\pi(f|s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))\cdot\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
@@
This batching setup is new in our application context compared to other atomic action approaches (e.g., Feng et al. (2021)), as our system-level overflow decisions are made on a discrete-time basis with multiple customers to be processed in a batch. The first advantage of this batching approach is that the atomic policy for the same class of customers remains consistent within the same decision epoch, which ensures a degree of fairness regardless of the pre-selected processing order $\boldsymbol{\sigma}(s)$. The second advantage is the improved time efficiency in both the simulation and policy update processes. Unlike the original sequential atomic decision setup, we do not need to update each atomic state or compute the state-dependent atomic action probabilities for each individual customer. The third advantage lies in the great simplification of calculating the objective function (11) for PPO. Specifically, recall that we need to compute the probability ratio
$r_{\theta,\eta}(f|s)=\frac{\pi_{\theta}(f|s)}{\pi_{\eta}(f|s)}$
at each decision epoch, where $\pi_{\theta}(f|s)$ and $\pi_{\eta}(f|s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively. Under the original setup, getting this ratio is computationally intensive because it involves summing over all possible atomic action sequences $\boldsymbol{a}$ corresponding to the system-level action $f$ for a given customer order $\boldsymbol{\sigma}(s)$ (i.e., the summation over $\boldsymbol{a}$ part in (5)). In contrast, under the batching setup, using (14) from Lemma 14, this ratio is greatly simplified to:
$r_{\theta,\eta}(f|s)=\prod_{i=1}^{J}\prod_{j=1}^{J}\left(\frac{\kappa_{\theta}(j|s,i)}{\kappa_{\eta}(j|s,i)}\right)^{f_{i,j}},$
@@
As in the original setup, it is possible that during implementation a sampled atomic action under the batching setup could become infeasible if the selected pool no longer has an available server or cannot accommodate the customer class. To handle this, we apply the same resampling technique to ensure a feasible action is selected. Other implementation details are as follows. We use Tensorflow v1 to build the policy network training pipeline and use the Ray package for parallel simulation data generation. The PPO algorithm iterates through three major modules: simulation data generation, policy evaluation, and policy improvement. For data generation, multiple actors (CPUs) run parallel simulations of equal length, and then the simulated data is concatenated into one long episode. For policy evaluation, we compute the linear value function approximations and the estimated advantage functions. For policy improvement, we update the policy network parameters $\theta$ using the Adam optimization method, running for $E$ training epochs (one training epoch corresponds to the number of complete passes through the entire training dataset). This process repeats until the average costs from two iterations differ by less than a small $\delta$. Table 1 summarizes important PPO hyper-parameters used in a ten-pool system in the case study in Section 6. More details on tuning these hyper-parameters are provided in Appendix 12.1.
@@ 6 Case Study
In this section, we evaluate the effectiveness of our tailored PPO algorithm through a case study on inpatient overflow assignment in large hospital systems, including settings with up to twenty patient classes and twenty wards (pools). In Section 6.1, we introduce the case study settings. In Section 6.2, we show that our PPO algorithm achieves similar performance to the ADP method from Dai and Shi (2019) – considered the “state-of-the-art” – while significantly reducing computational time in the five-pool baseline, the largest system handled by Dai and Shi (2019). In Section 6.3, we focus the analysis on ten- and twenty-pool systems, where our method significantly outperforms several benchmark policies. In Section 6.4, we examine the algorithm’s performance across different hyper-parameter settings and highlight the critical role of our atomic policy design and tailored components. Finally, in Section 6.5 we summarize key takeaways and offers practical recommendations for managing inpatient overflow.
@@ 6.1 Model Settings
The model settings for our case study are based on the five-pool system from Dai and Shi (2019), which is calibrated using aggregate patient-flow statistics from a large teaching hospital in Asia. The calibrated system has 315 servers (inpatient beds) and a daily arrival rate of 70 customers (patients). Time-varying arrival and discharge patterns are detailed in Appendix 12. Each class (medical specialty) has a dedicated primary ward, one preferred overflow ward, and two secondary overflow wards. The corresponding overflow costs are $B=(B_{1},B_{2})$, with $B_{1}<B_{2}$ to reflect ward preferences. We set the baseline values at $B=(30,35)$ for overflow costs and $C=6$ for holding costs. These cost parameters act as “tuning knobs,” which allow managers to adjust them to generate Pareto curves and balance congestion with overflow rates to meet target performance goals.
We extend this model to ten-pool and twenty-pool systems. The ten-pool system consists of two departments: a VIP department with smaller capacity and a regular department with larger capacity. Both departments follow the five-pool structure, and the total capacity is double that of the baseline five-pool system. Arrival rates are adjusted proportionally, as shown in Table 2. Overflow routes within each department remain unchanged (one preferred and two secondary overflow wards), but we also allow VIP patients to overflow to the regular department (not vice versa). Figure 3 provides a schematic representation of the feasible routes in the ten-pool system, using the GeMed (General Medicine) specialty as an example. The diagram illustrates both within-department routes (black solid lines and red/blue dashed lines) and cross-department routes (orange and green dotted lines), highlighting the overflow pathways available within and between departments. Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4})=(25,30,35,40)$, with lower costs for within-department overflows $(B_{1},B_{2})$ and higher costs for cross-department overflows $(B_{3},B_{4})$. The unit holding cost is set at $C=6$ for regular patients and $C=7$ for VIP patients. For the twenty-pool system, we assume it consists of two hospitals, each modeled as a ten-pool system with different capacities, as shown in Table 2. Patients can overflow within their hospital or to the other hospital. Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4},B_{5},B_{6},B_{7},B_{8})=(25,30,35,49,40,45,50,55)$, with lower costs for within-hospital overflows $(B_{1},\dots,B_{4})$ and higher costs for cross-hospital overflows $(B_{5},\dots,B_{8})$. The baseline unit holding costs remain the same, i.e., $C=6$ for regular and $C=7$ for VIP patients.
@@
Table 2: Parameter settings for five-pool, ten-pool and twenty-pool models.
@@ 6.2 Five-pool System: Comparing Performance with ADP
Table 3 compares the performance of our PPO algorithm with the ADP method in the five-pool model. We consider two setups: a balanced setting where each pool has equal capacity, and an unbalanced setting where some wards have less capacity and higher congestion. Each day is divided into eight decision epochs, and following Dai and Shi (2019), we fix the unit holding cost while varying the unit overflow cost. These experiments show that the PPO method achieves performance comparable to ADP (within 4% difference) – the best-performing method for the five-pool model to date – while significantly reducing computational time. Specifically, PPO completes one iteration in about two hours, compared to over ten hours for ADP, since its atomic policy design eliminates the need for exhaustive search over all feasible actions. Furthermore, in the simpler two-pool systems where the exact optimal policy is known, PPO achieves near-optimal results (within 3% difference). More results are available in Appendix 12.
@@
ADP 152.68$\pm$0.76 196.31$\pm$0.92 250.39$\pm$0.45 132.46$\pm$0.32 193.03$\pm$0.62 248.48$\pm$0.71
PPO 153.17$\pm$0.34 202.37$\pm$0.65 260.87$\pm$0.40 134.59$\pm$0.82 196.78$\pm$1.06 258.32$\pm$0.98
Table 3: Five-pool system. We set $N_{1}=60,N_{2}=64,N_{3}=67,N_{4}=N_{5}=62$ in unbalanced setting and $N_{1}=\cdots=N_{5}=63$ in balanced setting. Original atomic decision (non-batched) is used here.
@@ 6.3 Large Systems: Algorithm Scalability
In larger systems with more than five pools, the ADP method becomes infeasible due to the need to exhaustively search over the combinatorially large action space. Therefore, we compare the performance of the PPO algorithm against three benchmark policies introduced in Dai and Shi (2019): (a) the complete-overflow policy, where overflow assignments are made as many as possible at each decision epoch; (b) the midnight policy, allowing overflow assignments only at midnight each day; and (c) the empirical policy, which reflects current practice by allowing overflow assignments only during nighttime (7 PM to 7 AM the next day). Figure 4 shows the performance comparison in ten-pool and twenty-pool systems, where PPO significantly outperforms all three naive policies. In the ten-pool system, PPO achieves an average cost of 300, approximately 23% lower than the best-performing benchmark (the empirical policy with an average cost of 390). For the twenty-pool system, the improvement is similar: PPO reduces the average cost to 1000, about 25% lower than the empirical policy (which has an average cost of 1250). The computational time is about 3.5 hours per iteration for the ten-pool system and about 8 hours per iteration for the twenty-pool systems. Note these are training times - once the policy networks are trained, the execution of the optimal policy is almost instantaneous in real-time deployment. These results highlight the scalability and effectiveness of our PPO algorithm in large-scale systems.
@@ 6.4 Importance of Tailored Designs
In this section, we evaluate the performance of different hyper-parameter combinations in the ten-pool system, using the values from Table 1 as the baseline. While hyper-parameters for NNs play a role in algorithm performance as traditionally understood, we find that they are less critical than the design of the policy network, basis functions, and batching setup – the tailored elements introduced in this paper. These results highlight the importance of adapting a general-purpose algorithm to the specific problem domain for optimal performance and computational efficiency.
Policy network design. We compare the performance of three policy NN designs in the ten-pool system across various neural network widths, depths, and training settings; see Figures 5-6. The figures plot the percentage improvement in average cost achieved by the PPO policy (after 10 iterations) compared to the best-performing benchmark, the empirical policy, thus, a higher percentage indicates greater improvement (more desirable). We make the following observations.
@@
Figure 5: Trade-off between performance and training days and epochs. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
@@
Figure 6: Trade-off between performance and NNs’ depth and width. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy). A value of “0” in the depth or width indicates no hidden layer, which is possible for fully-connected or fully-separate networks but not for the partially-shared network due to its design structure.
Other tailored designs. Table 4 compares the average costs from using our proposed basis function against those used in Dai and Shi (2019), with varying amounts of simulation data. The table shows that our proposed basis functions are much less sensitive to the number of simulation days per actor. While the results with a larger dataset (10,000 days) are similar between the two sets of basis functions, the advantage of our proposed one becomes more evident with fewer simulation days (1,000 days) – the more practical choice for real-world implementation of the algorithm. Moreover, the results in Appendix 12.2 show that our proposed basis functions consistently obtains more accurate approximation among the top 25% most frequently visited states.
Batching is another key design for improving time efficiency, as discussed in Section 5.3. We implement the original atomic decision process following Feng et al. (2021) for our problem and compare it with our batching setup. The batching setup delivers similar performance to the original but reduces computational time by approximately 50% (e.g., from two hours to one hour); see Appendix 12.3. These findings show the value of our tailored designs in enhancing the time efficiency of the PPO algorithm. That is, by adapting to the queueing structure under the randomized atomic action setup, these tailored designs reduce sensitivity to the amount of simulation data. This not only speeds up the algorithm significantly but also makes it more practical for real-world applications where large-scale simulation data may be limited.
@@
Impact of other parameters. We present in Appendix 12.1 a detailed performance comparison under various combinations of hyper-parameters. The key takeaway remains consistent with our earlier discussion: our tailored PPO design captures the most salient features of the problem and is of first-order importance in reducing long-run average costs and improving computational efficiency. While tuning other hyper-parameters, such as the number of simulation days, training epochs, and policy update iterations, can further enhance performance, their impact is relatively marginal.
From the comparison results, we recommend increasing training epochs before expanding simulation data or iterations to maximize the use of existing data. Additionally, we identify two key techniques that further improve the algorithm’s efficiency:
  1. 1.
    Reuse of simulation data: Since PPO allows only small policy changes in each iteration, simulation data from a few previous iterations can be reused to reduce variance in estimating the value functions in the current iteration without introducing significant bias. This improves sample efficiency. Our experiments suggest that reusing data from the previous three iterations achieve the best results; see Table 12 in Appendix 12.3.
  2. 2.
    Adaptive clip size: A larger clip size can speed up policy improvement by allowing more substantial changes early in the training. However, it may hinder convergence to a stable policy later. Our experiments recommend starting with a larger clip size (e.g., 0.5) and reducing it as the policy update progresses; see Table 13 in Appendix 12.3.
6.5 Practical Implementation
A key strength of our approach is its domain adaptation: leveraging queueing structures and problem-specific characteristics to develop more effective and efficient algorithms. By integrating a tailored policy network, customized basis functions, and a batching setup, our PPO algorithm requires significantly less simulation data than typical reinforcement learning or ADP-based approaches. This not only accelerates training but also makes the algorithm more practical in environments with limited data or computational resources. The most important insights from our case study findings lies here as well, that is, the domain-aware adaptation is more critical to algorithm performance than fine-tuning neural network parameters in the original general-purposed version, which can be ad hoc and lack systematic guidance.

### v4

@@ 5.3 Batching Setup and Other Implementation Details
In the general atomic decision process, each of the atomic actions sequentially depends on the atomic states as illustrated in Figure 1. We find that decision ordering has limited impact on overall performance; see Appendix 12.2.4. Motivated by this observation, we propose a batching setup to simplify the decision process, which uses the same initial state $s=s^{1}$ for all atomic actions in a batch. Specifically, given $s$ and a chosen customer class order $\boldsymbol{\sigma}(s)$, the probability of conducting atomic action sequence a becomes $\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))$. In other words, during this batch, since the input state $s$ no longer changes, we prescribe atomic actions to the customers from the same class $i$ following the same distribution $\kappa(\cdot|s,i)$. Under the assumption that any system-level action $f$ satisfying $\sum_{j}f_{i,j}=q_{i},\forall i$ is feasible, we have the following lemma.
@@
$\pi(f\mid s)=\prod_{i=1}^{J}\frac{q_{i}!}{\prod_{j=1}^{J}f_{i,j}!}\prod_{j=1}^{J}\kappa(j|s,i)^{f_{i,j}}.$ (14)
@@
$\pi(f\mid s)=$ $\sum_{\textbf{a}}\prod_{n=1}^{q}\kappa(a^{n}|s,\sigma^{n}(s))\cdot\mathbbm{1}\left(f_{i,j}=\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i),i,j\in\mathcal{J}\right)$
@@
The first advantage of this batching approach is that the atomic policy for the same class of customers remains consistent within the same decision epoch, which ensures a degree of fairness regardless of the pre-selected processing order $\boldsymbol{\sigma}(s)$. The second advantage is the improved time efficiency in both the simulation and policy update processes. Unlike the original sequential atomic decision setup, we do not need to update each atomic state or compute the state-dependent atomic action probabilities for each individual customer. The third advantage lies in the great simplification of evaluating (11). Specifically, recall that we need to compute the probability ratio
$r_{\theta,\eta}(f\mid s)=\frac{\pi_{\theta}(f\mid s)}{\pi_{\eta}(f\mid s)}$
at each decision epoch, where $\pi_{\theta}(f\mid s)$ and $\pi_{\eta}(f\mid s)$ are defined as in (5), with $\kappa$ replaced by $\kappa_{\theta}$ and $\kappa_{\eta}$, respectively. Directly calculating this ratio is computationally intensive because it involves summing over all possible atomic action sequences $\boldsymbol{a}$ corresponding to the system-level action $f$ for a given customer order $\boldsymbol{\sigma}(s)$ (i.e., the summation over $\boldsymbol{a}$ part in (5)). In contrast, under the batching setup, using (14) from Lemma 14, this ratio is greatly simplified to:
$r_{\theta,\eta}(f\mid s)=\prod_{i=1}^{J}\prod_{j=1}^{J}\left(\frac{\kappa_{\theta}(j|s,i)}{\kappa_{\eta}(j|s,i)}\right)^{f_{i,j}},$ (15)
@@
As in the original setup, it is also possible that a sampled action under the batching setup could become infeasible if the selected pool no longer has an available server or cannot accommodate the customer class. To handle this, we apply the same resampling technique to ensure a feasible action is selected. Other implementation details are as follows. We use Tensorflow v1 to build the policy network training pipeline and use the Ray package for parallel simulation data generation. The PPO algorithm iterates through three major modules: simulation data generation, policy evaluation, and policy improvement. For data generation, multiple actors (CPUs) run parallel simulations of equal length, and then the simulated data is concatenated into one long trajectory, which constitutes the entire training dataset. For policy evaluation, we compute the linear value function approximations and the estimated advantage functions. For policy improvement, we update the policy network parameters $\theta$ using the Adam optimization method, running for $E$ training epochs (one training epoch corresponds to the number of complete passes through the entire training dataset). This process repeats until the average costs from two iterations differ by less than a small $\delta$. Table 1 summarizes important PPO hyper-parameters used in a ten-pool system in the case study in Section 6. More details on tuning these hyper-parameters are provided in Appendix 12.1.
@@ 6 Case Study
In this section, we evaluate the effectiveness of our tailored PPO algorithm through a case study on inpatient overflow assignment in large hospital systems, including settings with up to twenty patient classes and twenty wards (pools). In Section 6.1, we introduce the case study settings. In Section 6.2, we show that our PPO algorithm achieves similar performance to the ADP method from Dai and Shi (2019) – considered the “state-of-the-art” – while significantly reducing computational time in the five-pool baseline, the largest system handled by Dai and Shi (2019). In Section 6.3, we show the scalability in ten- and twenty-pool systems. In Section 6.4, we examine the algorithm’s performance across different hyper-parameter settings. We also conduct a detailed ablation study to highlight the critical role of our atomic policy design and the value of each tailored component. Finally, in Section 6.5 we showcase that the policy from the algorithm output is interpretable, and we summarize key takeaways for managing inpatient overflow.
@@ 6.1 Model Settings
The model settings for our case study are based on the five-pool system from Dai and Shi (2019), which is calibrated using aggregate patient-flow statistics from a large teaching hospital in Asia. The calibrated system has 315 servers (inpatient beds) and a daily arrival rate of 70 customers (patients). Time-varying arrival and discharge patterns are detailed in Appendix 12. Each class (medical specialty) has a dedicated primary ward, one preferred overflow ward, and two secondary overflow wards. The corresponding overflow costs are $B=(B_{1},B_{2})$, with $B_{1}<B_{2}$ to reflect ward preferences. We set the baseline values at $B=(30,35)$ for overflow costs and $C=6$ for holding costs. These cost parameters act as “tuning knobs,” which allow managers to adjust them to generate Pareto curves and balance congestion with overflow rates to meet target performance goals; see more in Appendix 12.3.
We extend this model to ten-pool and twenty-pool systems. The ten-pool system consists of two departments: a VIP department with smaller capacity and a regular department with larger capacity. Both departments follow the five-pool structure, and the total capacity is double that of the baseline five-pool system. Arrival rates are adjusted proportionally, as shown in Table 2. Overflow routes within each department remain unchanged (one preferred and two secondary overflow wards), but we also allow VIP patients to overflow to the regular department (not vice versa). Figure 3 provides a schematic representation of the feasible routes in the ten-pool system, using the GeMed (General Medicine) specialty as an example. The diagram illustrates both within-department routes (black solid lines and red/blue dashed lines) and cross-department routes (orange and green dotted lines), highlighting the overflow pathways available within and between departments. Overflow costs are set at $B=(B_{1},B_{2},B_{3},B_{4})=(25,30,35,40)$, with lower costs for within-department overflows $(B_{1},B_{2})$ and higher costs for cross-department overflows $(B_{3},B_{4})$. The unit holding cost is set at $C=6$ for regular patients and $C=7$ for VIP patients. For the twenty-pool system, we assume it consists of two hospitals, each modeled as a ten-pool system with different capacities, as shown in Table 2. Patients can overflow within their hospital or to the other hospital. Overflow costs are set at $B=(B_{1},\dots,B_{8})=(25,30,35,49,40,45,50,55)$, with lower costs for within-hospital overflows $(B_{1},\dots,B_{4})$ and higher costs for cross-hospital overflows $(B_{5},\dots,B_{8})$. The unit holding costs remain the same.
@@
Table 2: Parameter settings for five-pool, ten-pool and twenty-pool models; the latter two are formed by clustering and proportionally scaling the five-pool model in arrivals and capacities.
@@ 6.2 Five-pool System: Comparing Performance with ADP
Table 3 compares the performance of our PPO algorithm with the ADP method in the five-pool model. We consider two setups: a balanced setting where each pool has equal capacity, and an unbalanced setting where some wards have less capacity and higher congestion. Each day is divided into eight decision epochs, and following Dai and Shi (2019), we fix the unit holding cost while varying the unit overflow cost. These experiments show that the PPO method achieves performance comparable to ADP (within 4% difference) – the best-performing method for the five-pool model to date – while significantly reducing computational time. Specifically, PPO completes one iteration in about two hours, compared to over ten hours for ADP, since its atomic policy design eliminates the need for exhaustive search over all feasible actions. Furthermore, in the simpler two-pool systems where the exact optimal policy is known, PPO achieves near-optimal results (within 3% difference).
@@
[…]
Table 3: Five-pool system: long-run average costs. We set $N_{1}=60,N_{2}=64,N_{3}=67,N_{4}=N_{5}=62$ in unbalanced setting and $N_{1}=\cdots=N_{5}=63$ in balanced setting. Original atomic decision (non-batched) is used here.
@@ 6.3 Large Systems: Algorithm Scalability
In larger systems with more than five pools, the ADP method becomes infeasible due to the need to exhaustively search over the combinatorially large action space. Therefore, we compare the performance of the PPO algorithm against three benchmark policies introduced in Dai and Shi (2019): (a) the complete-overflow policy, where overflow assignments are allowed at each decision epoch; (b) the midnight policy, allowing overflow assignments only at midnight each day; and (c) the empirical policy, which reflects current practice by allowing overflow assignments only during nighttime (7 pm to 7 am the next day). In all three policies, when overflow assignments are allowed, waiting patients are assigned one by one: first to their preferred overflow ward if feasible, and then to a randomly chosen secondary ward, until no waiting patients remain or no idle beds are available. Figure 4 shows the performance comparison in ten and twenty-pool systems, where PPO significantly outperforms all benchmark policies. In the ten-pool system, PPO achieves an average cost of 300, approximately 23% lower than the best-performing benchmark (the empirical policy with an average cost of 390). For the twenty-pool system, the improvement is similar: PPO reduces the average cost to 1000, about 25% lower than the empirical policy (which has an average cost of 1250). The computational time is about 3.5 hours per iteration for the ten-pool system and about 8 hours per iteration for the twenty-pool systems. Note these are training times; once the policy networks are trained, the execution of the policy is almost instantaneous in real-time deployment. These results highlight the scalability and effectiveness of our PPO algorithm in large-scale systems.
@@ 6.4 Importance of Tailored Designs
In this section, we evaluate the performance of different hyper-parameter combinations in the ten-pool system, using the values from Table 1 as the baseline. While hyper-parameters for NNs play a role in algorithm performance as traditionally understood, we find that they are less critical than the design of the policy network, basis functions, and batching setup – the tailored elements introduced in this paper. These results highlight the importance of adapting a general-purpose algorithm to the specific problem domain for optimal performance and computational efficiency. Table 4 summarizes our ablation study and the relative impact of each tailored component on performance and computation time.
Configuration Performance Change Time Change
Full algorithm
Remove PPO (atomic + greedy search) $\downarrow\downarrow$ $\uparrow\uparrow$
Remove tailored policy network (alternative NN designs) $\downarrow\downarrow$ $\sim$
Remove tailored basis (basis in Dai and Shi (2019)) $\downarrow$ $\sim$
Remove LSTD (NN critic) $\downarrow$ $\uparrow$
Remove batching (sequential updates, various orders) $\sim$ $\uparrow\uparrow$
Fewer epochs $\downarrow$ $\downarrow$
Fewer simulation days $\sim$ $\downarrow$
Table 4: Ablation study summary. For each variant, the term in parentheses indicates the replacement mechanism. $\uparrow\uparrow$” or $\downarrow\downarrow$” indicate large increase or decrease in performance and computation time, compared with the full algorithm (integrating all elements: Atomic + PPO + Tailored Policy Network + Tailored LSTD + Batching); $\uparrow$” or $\downarrow$” moderate increase or decrease; and $\sim$” negligible change.
Is PPO necessary? Row 2 of Table 4 examines the effect of removing PPO and replacing it with a conventional value-based method (ADP) paired with atomic actions. In this setup, policy updates are performed via greedy search over all possible atomic actions. While atomic actions reduce the action space and make updates tractable, combining them with greedy search introduces significant instability due to the well-known “chattering” phenomenon (Bertsekas and others 2011). The average cost varying between 259.24 and 282.77, more than 27% increase from our PPO algorithm (see Appendix 12.2.2 for details). The chattering arises from the compounding effect of (i) the greedy policy updates (ii) the imperfect approximation of the value function, even with carefully chosen basis functions. These challenges are especially pronounced in long-run average cost settings, where small misestimations in value can lead to oscillatory behavior across iterations. Moreover, the atomic setup, which evaluates actions at an individual (rather than system) level, further exacerbates the chattering since each action’s marginal impact is highly sensitive to estimation noise. In contrast, pairing atomic actions with PPO mitigates these issues via incremental policy updates (as formalized in Proposition 1) and updating actions simultaneously via the system-level policy ratio (10). Together, these results highlight that the combination of atomic actions with PPO is essential in our infinite-horizon, long-run average context.
Policy network design. We compare the performance of three policy NN designs in the ten-pool system across various neural network widths, depths, and training settings; Row 3 of Table 4 summarizes the impact; Figures 5-6 show the detailed results. The figures plot the percentage improvement in average cost achieved by the PPO policy (after 10 iterations) compared to the best-performing benchmark, the empirical policy, thus, a higher percentage indicates greater improvement (more desirable). We make the following observations.
@@
Figure 5: Impact of training days and epochs. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy).
@@
Figure 6: Impact of NNs’ depth and width. Performance is measured as the % improvement in average cost between the PPO policy and the best-performing benchmark (the empirical policy). A value of “0” in the depth or width indicates no hidden layer, which is possible for fully-connected or fully-separate networks but not for the partially-shared network due to its design structure.
Other tailored designs. Rows 4-5 of Table 4 report the impact of replacing our tailored LSTD with (a) using basis functions in Dai and Shi (2019) and (b) using a NN-based value network, respectively. Table 5 further compares the average costs of using our proposed basis functions against those in Dai and Shi (2019), under different amounts of simulation data. The results show that our tailored basis functions are much less sensitive to the number of simulation days: while performance is similar with a large dataset (10,000 days), the advantage of our proposed functions becomes more evident with fewer simulation days (1,000 days), which is more practical for real-world implementation. Using the NN-based value network, on the other hand, shows worse final performance and requires extra time for training. This is likely due to higher variance and less stable value estimates for training the NN. See Appendix 12.2.3 for detailed comparison.
Row 6 of Table 4 compares the batching approximation and the original sequential atomic updates. The batching setup substantially reduces computational time ($\sim 50\%$) while maintaining similar performance. These findings show the value of our tailored designs in enhancing the time efficiency of the PPO algorithm. This not only speeds up the algorithm significantly but also makes it more practical for real-world applications where large-scale simulation data may be limited.
@@
We present in Appendix 12.1 a detailed performance comparison under various combinations of hyper-parameters (remaining rows of Table 4). The key takeaway remains consistent with our earlier discussion: our tailored PPO design captures the most salient features of the problem and is of first-order importance in reducing long-run average costs and improving computational efficiency. While tuning other hyper-parameters, such as the number of simulation days, training epochs, and policy update iterations, can further enhance performance, their impact is relatively marginal.
6.5 Practical Implementation and Policy Interpretability
In this section, we discuss how to implement our approach in practice, examine the policy interpretability since this is key for future implementation, and summarize actionable insights.